A man drives a car a distance of 200km at an average speed of 44km/hr what must be his average speed for the next 220km if he is to cover the total distance in 9 hours


1
Expert's answer
2022-03-20T18:47:58-0400

The total time of motion

"t=t_1+t_2=\frac{d_1}{v_1}+\frac{d_2}{v_2}"

Hence

"\frac{200}{44}+\frac{220}{v_2}=9""v_2=49.4\:\rm km/hr"